OpenStudy (anonymous):

Solve for x. 8x^2-5=11

OpenStudy (moonlitfate):

All right! Do you know what the first step to solving the equation would be at all?

OpenStudy (anonymous):

nope !:(

OpenStudy (moonlitfate):

Basically, what you're going to do is try isolate x (get it by itself) in order to solve for it.

OpenStudy (anonymous):

alright, and i get to but i dont know what kind of two they have weird answer choices for me?

OpenStudy (anonymous):

2*

OpenStudy (anonymous):

is it +-2 or the square root =-2?

OpenStudy (moonlitfate):

\[8x^2 - 5 = 11\] Add 5 to both sides of the equation to isolate x: \[8x^2-5(+5) = 11 +5\] The 5 on the left will cancel out leaving you with:\[8x^2 = 16\] \[x^2 = 16\] \[x = \sqrt{16}\] \[x = 4\]

OpenStudy (moonlitfate):

Oh wait, I skipped a step; sorry. x.x It should be....

OpenStudy (anonymous):

alright but all my choices have to do with a 2

OpenStudy (anonymous):

wait no they dont so its 4?

OpenStudy (moonlitfate):

No, I skipped a step; so ignore that for now.

OpenStudy (moonlitfate):

\[8x^2 = 16\] \[x^2 = 2\] \[x = \sqrt{2}\]

OpenStudy (moonlitfate):

Is it asking for two answers, though?

OpenStudy (anonymous):

nope just one thank you so much its the squear root of two?

OpenStudy (moonlitfate):

\[\pm \sqrt{2}\] ; you can't have a negative under a radical; since that would be a completely different concept of Algebra. :p

OpenStudy (anonymous):

what about this one: What constant term should be used to complete the square? x^2 - 5x + _____ = 7

OpenStudy (moonlitfate):

Let me see if I can remember this, been a while. :)

OpenStudy (moonlitfate):

\[x^2-5x+ ? = 7\] What you're going to take the middle term of your polynomial divide it by 2 and square it, if that makes sense.

OpenStudy (moonlitfate):

So the first step:

OpenStudy (moonlitfate):

\[x^2-5x+\frac{ 10 }{ 4 }=7\]

OpenStudy (anonymous):

alright ...got it

OpenStudy (moonlitfate):

Since you added 10/4 on the left, you're going to same on the right. \[x^2-5x+\frac{ 10 }{ 4 }=7+\frac{ 10 }{ 4 }\]

OpenStudy (moonlitfate):

x^2-5x+10/4 = 38/4 \[(x-\frac{ 5 }{ 2 })^2=\frac{ 38 }{ 4 }\]

OpenStudy (moonlitfate):

@exzaireah -- Is this helping at all? :x

OpenStudy (anonymous):

yes im understandin just not any of my answer choices.

OpenStudy (moonlitfate):

What are the choices given to you? Just to make sure I'm doing this correctly; I don't want to confuse you by doing the wrong thing.